Two spaceships A and B are exploring a new planet. Relative to this planet, spaceship A has a speed of 0.61c, and spaceship B has a speed of 0.72c. What is the ratio DA/DB of the values for the planet's diameter that each spaceship measures in a direction that is parallel to its motion?
Answer #1

Apply special throey of relativity equations.

DA = D*sqrt(1 - (vA/c)^2)

= D*sqrt(1 - (0.61*c/c)^2)

= D*sqrt(1 - 0.61^2)

= D*0.7924

simillary,
DB = D*sqrt(1 - (vB/c)^2)

= D*sqrt(1 - (0.72*c/c)^2)

= D*sqrt(1 - 0.72^2)

= D*0.6939

so, DA/DB = D*0.7924/(D*0.6939)

= 1.142